OpenStudy (surryyy):

In a square, the sum of the lengths of three sides is 48 and in a rectangle the difference between a length and a width is 7 in. If a width of the rectangle is the same as a side of the square, how long is a length of the rectangle? Answer 17 23 32 37

OpenStudy (perl):

Let S = side of square L = length of rectangle W = width of rectangle We are given 3S = 48 L- W = 7 W = s can you solve it now

OpenStudy (perl):

hint, first solve for S

OpenStudy (surryyy):

s = 48/3 s = 16 L = s + 7 L = 16 + 7 L = 23 is the length of the rectangle

OpenStudy (surryyy):

is it correct?